How much money would each person owe if the national debt were evenly distributed among all living people in the United States? I need to figure out what the individual share's would be for each situation. The second number is the population (the space in between the two numbers)
National Debt in Billions / U.S. Population in Millions / Individual Share
1.259 140
2.930 227
3.4,974 266
4.14,025 309
5. 18,100 320

Answer by stanbon(75887)

How much money would each person owe if the national debt were evenly distributed among all living people in the United States? I need to figure out what the individual share's would be for each situation. The second number is the population (the space in between the two numbers)
National Debt in Billions / U.S. Population in Millions / Individual Share
1. 259.....................140
Debt per person = 259*10^9/(140*10^6) = 1.85*10^3 = $1850 per person
2. 930.................... 227
Debt per person = 950*10^9/(225*10^6) = 4.22*10^3 = $4220 per person
--------
3.4,974 266
4.14,025 309
---------
5. 18,100 320
Ans:: 56.5625*10^3 = $56,562,500 per person
----------------------
Cheers,
Stan H.
---------